OpenStudy (anonymous):

What is the slope-intercept form of the equation y space minus space 7 space equals space minus 5 over 2 open parentheses x space plus space 4 close parentheses?

OpenStudy (welshfella):

y - 7 = -5/2(x + 4) you need to rearrange this to the form y = mx + b

OpenStudy (anonymous):

thats what i dont understand

OpenStudy (welshfella):

first add 7 to both sides then distribute the -5/2 over the parentheses finally simplify the right side

OpenStudy (anonymous):

ok hold on let me d that then see if im correct

OpenStudy (anonymous):

is it y=-5/2x - 3?

OpenStudy (welshfella):

right!

OpenStudy (anonymous):

omg thank you!!!!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

-5/2 is the slope of the line and -3 is the intercept on the y-axis

OpenStudy (welshfella):

though it is better to write it as y = (-5/2) x - 3 as it could be interpreted as -5 ------ 2x - 3

OpenStudy (anonymous):

okay thanks! can you help me with another question?

OpenStudy (welshfella):

ok 1 more

OpenStudy (anonymous):

Which is the equation for the line perpendicular to y space equals space minus 5 over 3 x space plus 11 1 third and containing P(-2, 3)?

OpenStudy (welshfella):

y = (-5/3) x + 11 1/3 the slope of this line is -5/3 so the slope of the line perpendicular to it = -1 / (-5/3) = 3/5

OpenStudy (anonymous):

so what do i do now?

OpenStudy (welshfella):

i'm a bit confused here - the line which is perpendicular contains the point P (-2,3) - right?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

ok so we need to use the general form of the point-slope equation y - y1 = m(x - x1) where m = slope (3/5) and (x1,y1) = (-2,3) plus these values into the above formula

OpenStudy (welshfella):

(if the slope of a line is m then the slope of a line perpendicular to it is -1/m)

OpenStudy (anonymous):

is it 3/5x + 4 1/5 ??

OpenStudy (welshfella):

y = (3/5) x + 4 1/5 Yes

OpenStudy (anonymous):

omg yay thanks so much!

OpenStudy (welshfella):

yw